Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-302_2000-308

Mission Operations Days: 2000/302 to 2000/308
October 28, 2000 00:00:00 GMT to November 03, 2000 00:00:00 GMT


  • MODIS is now operating via B/Redundant-Side Components (Except the FDDI Port)
  • SMIR Itwk/Vdet Values have been set to 79/110 (respectively)
  • All Band/Detector Gains have returned to their Initial Launch Values

Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)(New) Nominal
Calibration ElectronicsA Off; B On(New) Nominal
Control ProcessorA Off; B On(New) Nominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B On(New) Nominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 On(New) Nominal
Format ProcessorA Off; B On(New) Nominal
Power Supplies#1 Off; #2 On(New) Nominal
PV VIS FPAA Off; B On(New) Nominal
PV NIR FPAA Off; B On(New) Nominal
PV SMIR FPAA Off; B On(New) Nominal
PV LWIR FPAA Off; B On(New) Nominal
PC LWIR FPAA Off; B On(New) Nominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B On(New) Nominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B On(New) Nominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
(New) Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68(New) Nominal

This Week’s Completed MODIS Activities:

Saturday, October 28, 2000

None

Sunday, October 29, 2000

None


Transition to B-Side Sequence


Monday, October 30, 2000

2000/304 14:11:13 Real-Time Turn off MODIS proc: MOD_ALL_OFF
2000/304 14:12:50 Real-Time Turn on MODIS with PS2 and CPB proc: MOD_TURN_ON('PS2','CPB')
2000/304 14:15:35 Real-Time Disable TMON #68 commanding, FIFO double write proc: FS1_EXEC_TMON_GRP_COM(0,68,68)
2000/304 14:21:29 Real-Time Transition to Science Mode on the B-side proc: MOD_SCI_TXTN_B
2000/304 14:22:04 Real-Time Enable TMON #68 commanding, FIFO double write proc: FS1_EXEC_TMON_GRP_COM(1,68,68)
2000/304 14:24:43 Real-Time Reset Formatter to Upload Mode proc: MOD_FR_RESET_UPLD
2000/304 14:26:00 Real-Time Load Formatter Patch #1 proc: MOD_MEM_LOAD('AM1_MPR_MODIS_FR_ANOMALY2',2,'N','DISABLE',1,1)
2000/304 14:27:00 Real-Time Load Formatter Patch #2 proc: MOD_MEM_LOAD('AM1_MPR_MODIS_FR_RECOVERY',64,'N','DISABLE',1,2)
2000/304 14:27:35 Real-Time Restart Formatter with ASTATE = 0 proc: MOD_FR_RESTART(0)
2000/304 14:57:54 - 15:09:52 X-band Direct Broadcast service to Goddard (First look verifying collection of MODIS Science Data)
2000/304 15:01:00 Ground-Proc Set Yellow Limits for CP Status Word 6 (this produces a yellow limit if the Formatter Patch encounters an error not previously seen.) proc: MOD_ALARMS_CP_ST_06('ENABLE')
2000/304 15:02:00 Ground-Proc Set Yellow Limits for CP UART Reset (this produces a yellow limit if the CP loses communication with the Format Engine.) proc: MOD_ALARMS_URT_RST('ENABLE')


Tuesday, October 31, 2000

2000/305 19:12:36 Real-Time Set SMIR Itwk/Vdet to 79/110 proc: MOD_VDET_SWEEP('SMIR',0,0,79,110,0,0,1)
2000/305 20:21:52 ATC OA15: SDSM Open
2000/305 22:00:45 ATC OA16: SDSM Screened
2000/305 23:00:00 ATC Set Black Body Duty Cycle to 'FULL'
2000/305 23:00:30 ATC Set Black Body Temperature to '315K'

Wednesday, November 1, 2000

2000/306 06:00:00 ATC Set Black Body Duty Cycle to 'THIRD'
2000/306 06:00:30 ATC Set Black Body Temperature to '270K'

Thursday, November 2, 2000

2000/307 04:00:30 ATC Set Black Body Temperature to '290K'
2000/307 19:34 Real-Time Disable SRCA TMONs 66 and 67 processing As per B-Side Transition CAM discussion and submitted Terra Engineering Work Order. proc: FS1_EXEC_TMON_GRP_PROC(0,66,67)

Friday, November 3, 2000

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Friday, November 3, 2000

2000/308 21:44:49-22:42:20 ATC OA-23: SRCA Full Spatial

Saturday, November 4, 2000

2000/309 06:14:00-06:31:00 ATC Formatter Day Rate Science Collection during Nadir Ground Track Night. (3+ Granules)
2000/309 21:39:13-21:41:44 ATC OA-27: PV ECAL
2000/309 21:44:13-21:48:44 ATC OA-28: PC ECAL

Sunday, November 5, 2000

None

Monday, November 6, 2000

None

Tuesday, November 7, 2000

None

Wednesday, November 8, 2000

None

Thursday, November 9, 2000

None

Friday, November 10, 2000

None

Maneuvers:

Wednesday, November 15, 2000

The next MODIS Roll Maneuver is schedule for DOY 2000/320 22:41:59 to 22:58:59 for a roll angle of -11.5506 degrees.

MODIS Anomalies:

Continued monitoring for formatter errors and loss of cooler margin.

Attached to this report is a spreadsheet listing the times, addresses, formatter temperature, satellite ground track, and space craft day/night position for a select number of the formatter events that have occurred from DOY 2000/256 to DOY 2000/304. The final event count before the B-Side Transition was 2702. Due to the high number of recent events, not all details have been updated in the anomaly log.

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP). No new events have been witnessed by the FRFSP since the B-Side Transition on DOY 304.

MODIS Telemetry Trends:

The Radiative Cooler intermediate stage temperature has slowly been increasing since approximately DOY 2000/240. During this time the LWIR focal plane heater margin has been decreasing. The IOT is continuing to trend these telemetry mnemonics.

Non-MODIS Significant Events:

Wednesday, November 1, 2000

The Delta-V Maneuver was performed successfully.

No additional problems have been witnessed from the SFE-A since it was restarted on DOY 301.

As a reminder, the duration that the SFE-A was disabled: 2000/300 14:05:02 to 2000/301 14:06:19 (Just over 24 hours). NO Science Data will be available for this time period. All Science Data in this time frame has been permanently lost.

Limited Life Item Status:

SRCA 10W Lamp #1: 173.5 of 500 hours SRCA 10W Lamp #2: 135.9 of 500 hours SRCA 10W Lamp #3: 145.1 of 500 hours SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 556.1 of 4000 hours SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1453 of 3022 Movements Nadir Aperture Door: 532 of 1316 Movements Space View Door: 437 of 1316 Movements